# Beyond Diagonal RIS for 6G Non-Terrestrial Networks: Potentials and Challenges

### v1

@@ Abstract
Reconfigurable intelligent surface (RIS) has emerged as a promising technology in both terrestrial and non-terrestrial networks (NTNs) due to its ability to manipulate wireless environments for better connectivity. Significant studies have been focused on conventional RIS with diagonal phase response matrices. This simple RIS architecture, though less expensive, has limited flexibility in engineering the wireless channels. As the latest member of RIS technology, beyond diagonal RIS (BD-RIS) has recently been proposed in terrestrial setups. Due to the interconnected phase response elements, BD-RIS significantly enhances the control over the wireless environment. This work proposes the potential and challenges of BD-RIS in NTNs. We begin with the motivation and recent advances in BD-RIS. Subsequently, we discuss the fundamentals of BD-RIS and NTNs. We then outline the application of BD-RIS in NTNs, followed by a case study on BD-RIS enabled non-orthogonal multiple access low earth orbit satellite communication. Finally, we highlight challenges and research directions with concluding remarks.
@@ I Introduction
Fifth-generation wireless networks have primarily focused on optimizing transceiver designs to manage the uncontrollable wireless environment. However, as we move into the realm of sixth-generation (6G) and beyond networks, there is a shift towards new emerging concepts such as non-terrestrial networks (NTNs) and the coexistence of terrestrial and non-terrestrial networks. While this integration offers numerous benefits, it also presents the challenge of interference when these networks share the same spectrum resources [1]. To address this issue, one possible concept might be manipulating the wireless environment itself, thanks to the innovative technology of reconfigurable intelligent surfaces (RIS). RIS consists of reconfigurable elements capable of altering the propagation environment to enhance the channel gain of the received signals, thus enhancing the spectral efficiency of wireless communications. While the advantages of conventional diagonal RIS have been validated in various applications, it relies on a simple diagonal phase response matrix, where each element independently controls the phase of the incident signal without any cooperation with other elements [2]. This approach has some shortcomings: it only allows phase control, limiting passive beamforming capabilities, and it confines signal reflection to a single direction, thus limiting wireless coverage.
To address these limitations, a new member of RIS technology called beyond diagonal RIS (BD-RIS) has recently merged, which introduces connections among phase response elements that significantly enhance control over the reflected signals [3]. Moreover, the mathematical model for BD-RIS moves beyond the constraints of diagonal matrices, allowing for more sophisticated wave manipulations. Depending on how the phase response elements are connected, various BD-RIS architectures can be classified as single-connected, fully-connected, and group-connected BD-RIS, respectively. This leads to different characteristics of phase response matrices, i.e., block diagonal matrix for group-connected, unitary matrix for fully-connected, and diagonal for single-connected architecture. Note that single-connected is a special case of group-connected such that each group consists of a single phase response element. Fig. 1 envision a broad range of BD-RIS uses in 6G NTNs.
Figure 1: BD-RIS enabled 6G NTNs.
@@ I-A Related Works
Recently, some works have investigated the performance of BD-RIS in 6G wireless networks. In [4], Li et al. have proposed channel estimation and beamforming design for multi-antenna BD-RIS enabled wireless systems to investigate mean square error, spectral efficiency and sum rate. In [5], the authors have derived a closed-form solution for received channel gain maximization of BD-RIS enabled multi-user system by optimizing the active and passive beamforming. In [6], Esmaeilbeig et al. have adopted alternating optimization to maximize the signal-to-noise ratio in BD-RIS enabled integrated sensing and communication systems. In [7], Li et al. have computed the scaling law for received signal and maximized the sum rate of multi-user networks by designing an efficient algorithm for multi-sector BD-RIS. In [8], Mahmood et al. have optimized communication and computational resources to reduce the system latency in BD-RIS equipped unmanned aerial vehicle based mobile edge computing. In [9], Nerini et al. have derived a closed-form solution for designing efficient precoding at the transmitter, combining at the receiver, and beamforming at BD-RIS to maximize the average received sum power of the system. In [10], Soleymani et al. have optimized transmit beamforming and phase response to improve the performance of BD-RIS enabled multi-cell rate splitting multiple access networks. In [11], Liu et al. have jointly designed linear filter, transmit beamforming, and phase response to maximize the sum rate of BD-RIS enabled integrated sensing and communication. Moreover, Zhou et al. [12] have minimized the power consumption and maximized energy efficiency in BD-RIS enabled wireless networks by optimizing the transmit beamforming and BD-RIS phase response. In [13], the authors have jointly optimized the transmit precoder and phase response to maximize the sum rate of BD-RIS enabled multi-user systems. Table I provides the detailed summary of the above works on BD-RIS enabled 6G wireless networks.
Ref. Year Scenario Objective Findings Proposed solution
[4] 2024 Terrestrial CE and BF design Mean square error, SE, SR Used least square method for CE and efficient algorithm for beamforming design.
[5] 2024 Terrestrial Active and passive BF design SR, effective channel gain and time complexity Exploited two-stage low-complex joint active and passive beamforming at BS and BD-RIS.
[6] 2024 Terrestrial Phase response design for BD-RIS Signal-to-noise ratio Adopted an alternating optimization to jointly optimize the $\Phi$ for integrated sensing and communication.
[7] 2023 Terrestrial BF design for multi-sector BD-RIS SR Derived the scaling law for received signal and Designed an efficient algorithm for multi-sector BD-RIS.
[8] 2024 Terrestrial BD-RIS deployment, computational and communication resources Latency, max-min rate, task segmentation and resource allocation Proposed a joint iterative optimization approach for BD-RIS deployment, computational and communication resources.
[9] 2024 Terrestrial Precoding, combining, and BF design Average received sum power and computational complexity Derived a closed-form solution for precoding at transmitter, combining at receiver and BF at BD-RIS.
[10] 2024 Terrestrial BF and phase response design Fairness rate and fairness EE Proposed majorization minimization and alternating optimization to iteratively solve transmit BF and phase response problem for fairness rate and EE improvement.
[11] 2024 Terrestrial Linear filter, transmit BF and BD-RIS phase response design SR Used majorization-minimization and penalty-dual decomposition methods to solve the problem of joint linear filter, transmit BF, and phase response design.
[12] 2023 Terrestrial Transmit BF and BD-RIS phase response design SR, EE, power consumption Utilized a uni-fold optimization approach to solve the problem of total power consumption minimization and EE maximization.
[13] 2023 Terrestrial Transmit precoding and BD-RIS phase response design SR Employed fractional programming theory to solve the transmit precoding and phase response problem for sum rate maximization.
@@
TABLE I: Summary of different works on BD-IRS enabled 6G wireless networks.
Although some research has been done on BD-RIS enabled terrestrial networks, there is no existing literature that considers BD-RIS in NTNs. Motivated by this, we study the potential and challenges of BD-RIS in 6G NTNs. The article first discusses the fundamentals of NTNs and BD-RIS technology. Then, it explains the applications of BD-RIS in NTNs. Subsequently, it provides a case study on BD-RIS enabled low earth orbit (LEO) satellite communication using downlink non-orthogonal multiple access (NOMA). Finally, this work highlights challenges and future research directions in BD-RIS enabled 6G NTNs.
@@ II Non-Terrestrial Networks
NTNs is a sophisticated and diverse network that consists of satellite platforms and aerial platforms. It combines the wide coverage of satellite networks in the space layer and the adaptability of aerial networks in the atmospheric layer, as illustrated in Fig. 1.
@@ II-A Atmospheric Layer
The aerial network in the atmospheric layer primarily consists of UAVs and HAPSs [1]. The UAVs function at varying heights, spanning from hundreds to thousands of meters, and have the capability to cover distances of many kilometers. The operation is straightforward, the deployment is convenient, and the mobility is versatile. It has the capability to establish network connections and transmit signals through relays. However, the device’s operational duration varies from a few minutes to a few hours due to its limited battery capacity. On the other hand, HAPSs are situated in the stratosphere at an altitude ranging from 20 km to 50 km above the Earth’s surface. These platforms tend to remain in a fixed position. A network is established among many HAPSs by mmWave or THz connections. Because HAPSs are positioned at high altitudes, each HAPS has an extensive coverage area that can span up to 100 kilometers in circumference when the elevation angle is 10 degrees. Consequently, by reducing the number of HAPSs, it is possible to obtain a broader coverage area and expedite deployment. HAPS communication systems possess the attributes of cost-effectiveness, minimal latency, rapid deployment, and high capacity in contrast to satellite systems.
The aerial network possesses exceptional flexibility and scalability, making it suitable for a wide range of applications such as emergency communication, network traffic offloading, and temporary communication. For instance, in the event of natural disasters causing the destruction of the terrestrial network, the aerial network can be swiftly deployed to restore connection. Furthermore, an aerial network exhibits a reduced expenditure and decreased delay compared to a satellite network. Nevertheless, the limitations on energy consumption and weight have resulted in a significant drawback for the aerial network, namely its limited endurance period. Moreover, due to the HAPSs restricted height, it can easily cause significant interference to the terrestrial network.
@@ II-B Space Layer
This layer includes different satellite networks. Based on height, the satellite network has three tiers: GEO, MEO, and LEO [1]. GEO satellites form the core space-based network, whereas MEO and LEO satellites form the access network that connects to the aerial or terrestrial network. GEO satellites are set at 35,786 kilometers and the transmission require 0.27 seconds to reach from GEO to ground terminal. GEO satellites cover up to one-third of Earth’s surface. Except for the South and North Poles, three GEO satellites evenly spaced offer global communication. The majority of MEO satellites operate at 8000 miles and they require 0.1 second for signal transmission to the ground terminal. These satellites mostly deliver high-speed data and messages. Many satellites operate exclusively in this orbit for telecommunications and military use. LEO satellites orbit 500–1000 miles above sea level. From this orbit, the satellite’s signal reaches the ground station in 0.05 seconds. This track is best for low latency applications.
@@
III Modes and Advantages of BD-RIS in NTNs
@@
In this section, we first study the different modes BD-RIS technology and then we discuss its advantages compared to the conventional diagonal RIS technology.
III-A Modes of BD-RIS in NTNs
@@
BD-RIS can be classified into three modes, including reflective BD-RIS, hybrid BD-RIS, and multi-sector BD-RIS [13].
III-A1 Reflective BD-RIS

### v2

@@ Abstract
Reconfigurable intelligent surface (RIS) has emerged as a promising technology in both terrestrial and non-terrestrial networks (NTNs) due to its ability to manipulate wireless environments for better connectivity. Significant studies have been focused on conventional RIS with diagonal phase response matrices. This simple RIS architecture, though less expensive, has limited flexibility in engineering the wireless channels. As the latest member of RIS technology, beyond diagonal RIS (BD-RIS) has recently been proposed in terrestrial setups. Due to the interconnected phase response elements (PREs), BD-RIS significantly enhances the control over the wireless environment. This work proposes the potential and challenges of BD-RIS in NTNs. We begin with the motivation and recent advances in BD-RIS. Subsequently, we discuss the fundamentals of BD-RIS and NTNs. We then outline the application of BD-RIS in NTNs, followed by a case study on BD-RIS enabled non-orthogonal multiple access low earth orbit satellite communication. Finally, we highlight challenges and research directions with concluding remarks.
@@ I Introduction
Fifth-generation wireless networks have primarily focused on optimizing transceiver designs to manage the uncontrollable wireless environment. However, as we move into the realm of sixth-generation (6G) and beyond networks, there is a shift towards new emerging concepts such as non-terrestrial networks (NTNs) and the coexistence of terrestrial and non-terrestrial networks. While this integration offers numerous benefits, it also presents the challenge of interference when these networks share the same spectrum resources [1]. To address this issue, one possible concept might be manipulating the wireless environment itself, thanks to the innovative technology of reconfigurable intelligent surfaces (RIS). RIS consists of reconfigurable elements capable of altering the propagation environment to enhance the channel gain of the received signals, thus enhancing the spectral efficiency of wireless communications. While the advantages of conventional diagonal RIS (CD-RIS) have been validated in various applications, it relies on a simple diagonal phase response matrix (PRM), where each element independently controls the phase of the incident signal without any cooperation with other elements [2]. This approach has some shortcomings: it only allows phase control, limiting passive beamforming (BF) capabilities, and it confines signal reflection to a single direction, thus limiting wireless coverage.
To address these limitations, a new member of RIS technology called beyond diagonal RIS (BD-RIS) has recently merged, which introduces connections among phase response elements (PREs) that significantly enhance control over the reflected signals [3]. Moreover, the mathematical model for BD-RIS moves beyond the constraints of diagonal matrices, allowing for more sophisticated wave manipulations. Depending on how the PREs are connected, various BD-RIS architectures can be classified as single-connected, fully-connected, and group-connected BD-RIS, respectively. This leads to different characteristics of PRMs, i.e., block diagonal matrix for group-connected, unitary matrix for fully-connected, and diagonal for single-connected architecture. Note that single-connected is a special case of group-connected such that each group consists of a single PRE. Fig. 1 envision a broad range of BD-RIS uses in 6G NTNs. As can be seen in the figure, BD-RIS can be used as transmissive, reflecting, and hybrid modes. While it is true that reflective and hybrid BD-RIS are typically deployed to enhance communication when the line-of-sight (LoS) link is blocked, transmissive BD-RIS can be equipped with non-terrestrial platforms to actively control and optimize the signal propagation, even when the LoS is available [^1] The transmissive BD-RIS offers more advantages over the reflective BD-RIS for two main reasons. First, the reflective BD-RIS has self-interference because the feed antenna (transmitter) and receiver are on the same side, causing interference between the transmitted and reflected signals. On the other hand, the transmissive BD-RIS has the feed antenna and receivers on the opposite side, so there is no self-interference. Second, unlike reflective BD-RIS, we can design transmissive BD-RIS with enhanced operational bandwidth and aperture efficiency.. By doing this, the signals can be dynamically shaped and steered not only to improve coverage in areas where the LoS is obstructed but also to enhance overall signal strength and quality in various environments. This approach allows for more flexible and efficient use of non-terrestrial resources, particularly in scenarios where ground-based BD-RIS deployment alone might not address all the communication challenges. Moreover, one of the main benefits of transmissive BD-RIS is performs as a transmitter without necessitating complex signal processing, setting it apart from conventional large-antenna array systems that rely on complicated and energy-hunger RF modules, thereby incurring high hardware costs.
Figure 1: BD-RIS enabled 6G NTNs which consist of three layers, i.e., space layer, atmospheric layer, and earth surface.
@@ I-A Related Works
Recently, some works have investigated the performance of BD-RIS in 6G wireless networks. The research work of [3] have optimized phase response of BD-RIS in multiple input multiple output system to enhance the received signal gain. In [4], Li et al. have proposed channel estimation and BF design for multi-antenna BD-RIS aided wireless systems to investigate mean square error, spectral efficiency and sum rate. In [5], the authors have derived a closed-form solution for received channel gain maximization of BD-RIS aided multi-user system by optimizing the active and passive BF. In [6], Esmaeilbeig et al. have adopted alternating optimization to maximize the signal-to-noise ratio in BD-RIS aided integrated sensing and communication systems. In [7], Li et al. have computed the scaling law for received signal and maximized the sum rate of multi-user networks by designing an efficient algorithm for multi-sector BD-RIS. In [8], Mahmood et al. have optimized communication and computational resources to reduce the system latency in BD-RIS equipped unmanned aerial vehicle based mobile edge computing. In [9], Nerini et al. have derived a closed-form solution for designing efficient precoding at the transmitter, combining at the receiver, and BF at BD-RIS to maximize the average received sum power of the system. In [10], Soleymani et al. have optimized transmit BF and phase response (PR) to improve the performance of BD-RIS aided multi-cell rate splitting multiple access networks. In [11], Liu et al. have jointly designed linear filter, transmit BF, and PR to maximize the sum rate of BD-RIS aided integrated sensing and communication. Moreover, Zhou et al. [12] have minimized the power consumption and maximized energy efficiency in BD-RIS aided wireless networks by optimizing the transmit BF and BD-RIS PR. In [13], the authors have jointly optimized the transmit precoder and PR to maximize the sum rate of BD-RIS aided multi-user systems. The detailed comparison of the existing works is summarized in Table I.
Reference Year NTNs Objective Findings Proposed solution
[…]
@@
TABLE I: Summary of different works on BD-RIS aided 6G wireless networks.
Although some research has been done on BD-RIS aided terrestrial networks, there is no existing literature that considers BD-RIS in NTNs. Motivated by this, we study the potential and challenges of BD-RIS in 6G NTNs. The article first discusses the fundamentals of NTNs and BD-RIS technology. Then, it explains the applications of BD-RIS in NTNs. Subsequently, it provides a case study on BD-RIS aided LEO satellite communication using downlink non-orthogonal multiple access (NOMA). Finally, this work highlights challenges and future research directions in BD-RIS aided 6G NTNs.
@@ II Non-Terrestrial Networks
NTNs is a sophisticated and diverse network that consists of satellite platforms and aerial platforms and ground terminals. It combines the wide coverage of satellite networks in the space layer and the adaptability of aerial networks in the atmospheric layer, as illustrated in Fig. 1.
@@ II-A Atmospheric Layer
The aerial network in the atmospheric layer primarily consists of unmanned aerial vehicles (UAVs) and high altitude platform stations (HAPSs) [1]. The UAVs function at varying heights, spanning from hundreds to thousands of meters, and have the capability to cover distances of many kilometers (km). The operation is straightforward, the deployment is convenient, and the mobility is versatile. It has the capability to establish network connections and transmit signals through relays. However, the device’s operational duration varies from a few minutes to a few hours due to its limited battery capacity. On the other hand, HAPSs are situated in the stratosphere at an altitude ranging from 20 km to 50 km above the Earth’s surface. These platforms tend to remain in a fixed position. A network is established among many HAPSs by mmWave or THz connections. Because HAPSs are positioned at high altitudes, each HAPS has an extensive coverage area that can span up to 100 km in circumference when the elevation angle is 10 degrees. Consequently, by reducing the number of HAPSs, it is possible to obtain a broader coverage area and expedite deployment. HAPS communication systems possess the attributes of cost-effectiveness, minimal latency, rapid deployment, and high capacity in contrast to satellite systems.
Figure 2: Different architecture of BD-RIS: (a) Cell-wise single-connected BD-RIS architecture with 2 cells; (b) Cell-wise fully-connected BD-RIS architecture with 4 cells and 2 groups; (c) Element-wise group-connected BD-RIS with 2 groups.
The aerial network possesses exceptional flexibility and scalability, making it suitable for a wide range of applications such as emergency communication, network traffic offloading, and temporary communication. For instance, in the event of natural disasters causing the destruction of the terrestrial network, the aerial network can be swiftly deployed to restore the connection. Furthermore, an aerial network exhibits a reduced expenditure and decreased delay compared to a satellite network. Nevertheless, the limitations on energy consumption and weight have resulted in a significant drawback for the aerial network, namely its limited endurance period. Moreover, due to the HAPSs restricted height, it can easily cause significant interference to the terrestrial network.
@@ II-B Space Layer
This layer includes different satellite networks. Based on height, the satellite network has three tiers: geostationary orbit (GEO), medium Earth orbit (MEO), and LEO [1]. GEO satellites form the core space-based network, whereas MEO and LEO satellites form the access network that connects to the aerial or terrestrial network. GEO satellites are set at 35,786 km and the transmission requires 0.27 seconds to reach from GEO to the ground terminal. GEO satellites cover up to one-third of Earth’s surface. Except for the South and North Poles, three GEO satellites evenly spaced offer global communication. The majority of MEO satellites operate between 8,000 and 20,000 km altitude and they require 0.1 second for signal transmission to the ground terminal. These satellites mostly deliver high-speed data and messages. Many satellites operate exclusively in this orbit for telecommunications and military use. LEO satellites orbit between 500 and 2000 km above sea level. From this orbit, the satellite’s signal reaches the ground station in 0.05 seconds. This track is best for low-latency applications.
@@
II-C Channel Modeling in NTNs and Integration with BD-RIS
The channel modeling and characteristics in NTNs differ from those in terrestrial networks. For example, compared to terrestrial network channels, the channels in NTNs have challenges of high mobility, high path loss due to the large distance between satellites and ground terminals, Doppler shift caused by the satellite spinning, and signal blockage due to geographical obstructions and atmospheric absorptions. Besides that, the link budget in NTN systems is often constrained, especially for ground terminals with small antennas operating at higher frequencies, where signal attenuation is more prominent. BD-RIS can add more degrees of freedom in controlling and optimizing signal reflections, hence its integration with NTNs fits well with these channel properties. In particular, BD-RIS can dynamically change its phase shifts and reflection characteristics to match the fast-changing NTN channels, therefore adjusting to Doppler effects, signal fading, and multipath propagation. Furthermore, by concentrating signal reflections more accurately onto the user’s antenna, BD-RIS can improve the link budget by offsetting the great path loss and guaranteeing improved signal strength at the receiver. BD-RIS is a necessary instrument for maximizing NTN channel performance, especially in demanding link budget situations, since it allows one to enhance the effective signal gain.
Figure 3: Different modes BD-RIS supported by the same circuitry of reconfigurable impedance network: (a) Reflective BD-RIS, (b) Element wise reflective BD-RIS with group-connected architecture with 2 groups, where the group size is 4, (c) Transmissive BD-RIS, (d) Hybrid BD-RIS; (e) Cell wise group-connected BD-RIS architecture with 2 groups and 4 cells, where the size of each cell is 2; (f) Multi-sector BD-RIS; and (g) Cell wise single-connected multi-sector BD-RIS architecture with 3 sectors and 2 cells, where the cell size is 3.
III Preliminaries of BD-RIS
@@
In this section, we first study the hardware architectures of BD-RIS, followed by different operating modes of this technology, and then we also discuss its advantages compared to the CD-RIS technology.
III-A Hardware Architectures of BD-RIS
The hardware architectures of BD-RIS can be classified into three main categories: 1) single-connected BD-RIS, 2) fully-connected BD-RIS, and 3) group-connected BD-RIS, as illustrated in Fig 2. In the following, we briefly discuss each category principle and the matrices they produce.
III-A1 Single-connected BD-RIS
The hardware architecture of BD-RIS in this category consists of independent reconfigurable elements. The PREs on the surface of a single-connected BD-RIS are not interconnected and each element independently manipulates the incident electromagnetic wave [13]. A cell-wise [^2] In this work, we use the terms ‘element’ and ‘cell’ interchangeably based on the mode of the BD-RIS. Please note that we use ‘element” when we refer to BD-RIS operating in reflecting mode and use ‘cell when referring to BD-RIS operating in hybrid/multi-sector modes, where more than 1 element with uni-directional radiating patterns are back-to-back placed to form a cell. single-connected hybrid BD-RIS architecture with 4 cells, where each cell contains 2 elements, is illustrated in Fig. 2(a). A hybrid BD-RIS with $K$ PREs will provide diagonal PRMs as ${\bf\Phi}_{r}=\text{diag}(\phi_{r,1},\phi_{r,2}\dots\phi_{r,K})$ and ${\bf\Phi}_{t}=\text{diag}(\phi_{t,1},\phi_{t,2}\dots\phi_{t,K})$, where ${\bf\Phi}_{r}$ is the reflecting PRM and ${\bf\Phi}_{t}$ is the transmissive PRM. Furthermore, the PREs in each cell must satisfy as $|\phi_{r,k}|^{2}+|\phi_{t,k}|^{2}=1$, $\forall k\in K$.
III-A2 Fully-connected BD-RIS
Unlike single-connected BD-RIS, all the PREs in this architecture are interconnected through an impedance network. A cell-wise fully-connected hybrid BD-RIS architecture with 2 cells, where each cell consists of 2 PREs, is shown in Fig. 2(b). The PRMs ${\bf\Phi}_{r}$ and ${\bf\Phi}_{t}$ in this architecture are non-diagonal (full scattering matrices), resulting in improved system performance. The PRMs of this architecture satisfies as ${\bf\Phi}_{r}^{H}{\bf\Phi}_{r}+{\bf\Phi}_{t}^{H}{\bf\Phi}_{t}={\bf I}_{K}$, where ${\bf I}_{K}$ is an identity matrix.
III-A3 Group-connected BD-RIS
To reduce the complexity of fully-connected DB-RIS architecture when the number of PREs is very large, group-connected architecture is introduced, which divides PREs into multiple groups. Figure 2(b) illustrates an element-wise group-connected reflective BD-RIS with 2 groups, where each group consists of 2 elements with a fully-connected configuration. In this architecture, each group is fully-connected and all groups constitute a block diagonal matrix such as ${\bf\Phi}_{r}=\text{blkdiag}({\bf\Phi}_{r,1},{\bf\Phi}_{r,2},\dots,{\bf\Phi}_{r,G})$ with $G$ is the number of groups. If $\bar{K}=K/G$ denotes the number of elements in each group, then the PRM of each group satisfies ${\bf\Phi}^{H}_{r,g}{\bf\Phi}_{r,g}={\bf I}_{\bar{K}}$ with ${\bf\Phi}_{r,g}\in\mathcal{C}^{\bar{K}\times\bar{K}}$.
III-B Operating Modes of BD-RIS in NTNs
@@
Each architecture of BD-RIS can operate in three modes, including reflective BD-RIS, hybrid BD-RIS, and multi-sector BD-RIS.
III-B1 Reflective BD-RIS